# Promptus: Can Prompts Streaming Replace Video Streaming with Stable Diffusion

### v1

@@ 5.5 Performance on real-world traces
This section demonstrates the performance of Promptus under real network traces. We use the traces and testbed described in §5.1 to run Promptus and two baselines. Figure 13 shows the CDF and mean of the frame quality. Since the total length of the traces is greater than the total length of the test videos, we loop the test videos to run through the entire traces.
@@
Figure 13: The performance of Promptus under real network traces. On one hand, Promptus’s quality is overall higher than the baselines. On the other hand, Promptus can significantly reduce the ratio of severely distorted frames.
@@
First, Promptus’s quality is overall higher than the baselines. For example, the mean LPIPS of Promptus is 0.111 and 0.092 lower than VAE and H.265, respectively. Second, Promptus can significantly reduce the ratio of severely distorted frames. For instance, only 5.2% of frames in Promptus have LPIPS higher than 0.32, while VAE and H.265 have 94.5% and 96.9%, respectively. These improvements are partly due to Promptus’s excellent compression efficiency, enabling it to provide higher perceptual quality at the same bitrate. On the other hand, since Promptus sends raw prompts without encoding (such as entropy coding), it can precisely control the target, thus making full use of bandwidth.
@@ 5.6 Overhead
In this section, we analyze the overhead of Promptus’s generation. We conduct tests on an Nvidia 4090D GPU, using CUDA to accelerate Promptus. The resolution of the generated video is 512*512, and the rank of the prompt is 8.
Table 2: Overhead of generating a frame
Steps Time (ms) Memory (MB)
Prompt dequantization 0.016 -
Prompt composition 0.025 -
Prompt interpolation 0.013 -
Noised previous frame 0.012 -
Stable Diffusion generation 6.160 -
Total 6.226 8952
Table 2 shows the fine-grained overhead of each step in Promptus’s image generation. Specifically, from receiving a prompt to generating a frame, Promptus includes the following steps: prompt dequantization, prompt composition, prompt interpolation, adding noise to the previous frame, and Stable Diffusion image generation. Among them, most steps only involve simple linear computations, so the time overhead is almost negligible. In contrast, Stable Diffusion image generation accounts for the vast majority of the time overhead. Therefore, we adopt StreamDiffusion [28], which accelerates Stable Diffusion to real-time through batch processing techniques and implementation optimizations. The difference is that compared to the original StreamDiffusion, Promptus removes the CLIP module that converts text to embeddings, thus further accelerating image generation to 6.160 ms. In summary, the total time overhead of Promptus’s image generation is 6.226 ms, achieving real-time.
@@ 6 Limitation
The time overhead of prompt fitting. Although Promptus can achieve real-time video generation at the receiver, prompt fitting cannot be performed in real-time at the sender. This is because prompt fitting requires iterative gradient descent. Although a single iteration is fast (on the same order of magnitude as the time overhead of real-time generation), the total time overhead is high due to the large number of iterations required for convergence (such as 500 iterations). To address this, using more efficient gradient descent algorithms to reduce the number of iterations required for convergence is expected to accelerate prompt fitting to real-time.
The latency of prompt interpolation. At the receiver side, Promptus obtains the prompts of intermediate frames through prompt interpolation of keyframes. This means that if an intermediate frame needs to be generated and played, it is necessary to wait until the subsequent keyframe is received, which introduces additional latency. To address this issue, designing keyframe extrapolation algorithms to replace interpolation is a future research direction.
Non-uniform keyframes. According to §4.1, Promptus sends keyframes uniformly based on the keyframe interval. However, different segments of a video often have different rates of change. For rapidly changing segments, the distance between keyframes in the prompt space is large, resulting in a poor approximation of intermediate frames using linear interpolation. A solution is to reduce the keyframe interval and send keyframes more densely. However, there also exist smoothly changing segments in the video, where densely sending keyframes brings little improvement to quality, thus wasting bandwidth. In summary, sending keyframes uniformly in this paper is inefficient. In the future, adaptive keyframe needs to be designed to transmit densely in rapidly changing segments and sparsely in smoothly changing segments.
@@ 7 Conclusion
In this paper, we propose Promptus, a novel system that replaces video streaming with prompt streaming by inverting video frames into prompts for Stable Diffusion. To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed. To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced. For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed. Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS. By pioneering the replacement of video codecs with prompt inversion and introducing prompt streaming, Promptus revolutionizes efficient video communication, surpassing the limitations of traditional approaches and paving the way for a new era of video streaming.

### v2

@@ 5.7 Performance on real-world traces
Figure 16: The performance of Promptus under real network traces. On one hand, Promptus’s quality is overall higher than the baselines. On the other hand, Promptus can significantly reduce the ratio of severely distorted frames (whose LPIPS is higher than 0.32).
This section demonstrates the performance of Promptus under real-world network traces. Since the total length of the network traces does not match the total length of the test videos, we loop the videos to align with the traces.
@@
Figure 16 shows the CDF and mean of the frame quality. First, Promptus’s quality is overall higher than the baselines. For example, the mean LPIPS of Promptus is 0.111, 0.092, 0.047 and 0.063 lower than VAE, H.265, VQGAN and NAS, respectively. Second, Promptus can significantly reduce the ratio of severely distorted frames. For instance, only 5.2% of frames in Promptus have LPIPS higher than 0.32, while VAE, H.265, VQGAN and NAS have 94.5%, 96.9%, 53.2% and 98.2% respectively. These improvements are partly due to Promptus’s excellent transmission efficiency, enabling it to provide higher perceptual quality at the same bitrate. On the other hand, since Promptus sends raw prompts without entropy coding, it can precisely control the target bitrate, thus making full use of bandwidth.
5.8 Ablation study of interpolation methods and generators
Figure 17: Comparison of different interpolation methods. Each method is based on frame 0 and frame 10 to interpolate frames 1 to 9. The results demonstrate that latent interpolation fails to preserve the motion between frames, resulting in spatial overlaps and ghosting. This is because the adjacent frames in latent space are not linearly continuous, making the interpolation unreasonable. While pixel interpolation (huang2022rife, ) can preserve motion, it inevitably leads to noticeable artifacts in cases of occlusion, edges, or newly appearing objects due to incorrect matching, as illustrated by the green and red boxes in the figure. Our prompt interpolation successfully preserves motion while avoiding artifacts and keeping the edges sharp. Details can be found in §5.8
In this section, we compare different interpolation methods and analyze why it should be prompt interpolation rather than latent or pixel interpolation. We also demonstrate why it is necessary to use the complete Stable Diffusion as the generator for inversion rather than just the VAE (Decoder). Specifically, for latent interpolation, we remove the U-Net from the inversion framework (described in §3), retaining only the VAE Decoder. Thus, we randomly initialize the frames in the latent space and apply gradient descent for fitting. After finishing the fitting, we perform linear interpolation on the fitting results (latent variables) of the key frames to approximate the intermediate frames, just like in prompt interpolation. The results are shown in "End-to-end latent interpolation" in Figure 17. For pixel interpolation, we apply RIFE (huang2022rife, ) (a real-time video frame interpolation algorithm) to the ground truth images of the key frames, directly estimating the intermediate frames. The results are shown in "End-to-end pixel interpolation" in Figure 17. Additionally, to compare with Promptus based solely on interpolation methods, we also conduct experiments where key frames (both the latent variables and images) come from Promptus, represented as "Latent interpolation" and "Pixel interpolation" shown in Figure 17. Each of the above methods is based on frame 0 and frame 10 to interpolate frames 1 to 9.
First, the results demonstrate that latent interpolation fails to preserve the motion between frames, resulting in spatial overlaps and ghosting, as shown in the second and third rows of the Figure 17. This is because the adjacent frames in latent space are not linearly continuous, making the interpolation between frames unreasonable. This indicates that the VAE alone cannot serve as the generator for video inversion, since inter-frame compression does not work, and frame-by-frame transmission would result in significant bandwidth wastage. A feasible solution is to use codecs to encode the latent variables of each frame. However, as shown in Figure 11 and Figure 10, this solution performs worse than prompt interpolation.
Second, while pixel interpolation can preserve motion, it inevitably leads to noticeable artifacts in cases of occlusion, edges, or newly appearing objects due to incorrect matching, as illustrated by the green and red boxes in the Figure 17.
Third, our prompt interpolation successfully preserves motion while avoiding artifacts and keeping the edges sharp. This is because the U-Net converts frames from latent space to prompt space, where the frames are linearly continuous, allowing the interpolation to approximate motion. It proves that the U-Net (diffusion process) plays a significant role in compression performance.
@@ 5.9 Overhead
Figure 18: Visualization of fitting results at different iterations. The results indicate that for the first frame of a new video (or new scene), the fitting process takes thousands of iterations to converge since it starts from scratch. For subsequent frames, thanks to the continuous prompt space (§3.3) providing sufficient scene priors, the fitting process can converge in only a few hundred iterations. For example, the second frame in the figure requires only 120 iterations to achieve a visual quality comparable to that of 2000 iterations for the first frame.
In this section, we analyze the overhead of Promptus. We conduct tests on an Nvidia 4090D GPU, using CUDA to accelerate Promptus. The resolution of the generated video is 512*512, and the rank of the prompt is 8. The Stable Diffusion model has a total of 867M parameters. With more parameters (such as SDXL Turbo (sdxlturbo, )), the Stable Diffusion model has stronger generative ability, making prompt fitting easier and allowing for a higher compression ratio. However, this also leads to increased overhead in terms of memory usage and run time. Thus, the current version of Promptus employs SD 2.1 Turbo, which is relatively lightweight.
Table 1: Time overhead of generating a frame §5.9.
Steps Time (ms)
Dequantization 0.016
Prompt composition 0.025
Prompt interpolation 0.013
Noised frame 0.012
Stable Diffusion generation 6.160
Total 6.226
Generation overhead. Table 1 shows the fine-grained overhead of each step in Promptus’s image generation. Specifically, from receiving a prompt to generating a frame, Promptus includes the following steps: prompt dequantization, prompt composition, prompt interpolation, adding noise to the previous frame, and Stable Diffusion generation. Among them, most steps only involve simple linear computations, so the time overhead is almost negligible. In contrast, Stable Diffusion generation accounts for the vast majority of the time overhead. Therefore, following StreamDiffusion (kodaira2023streamdiffusion, ), we use TAESD and TensorRT to accelerate the Stable Diffusion to 6.16 ms. In summary, the total time overhead of Promptus’s image generation at the receiver is 6.226 ms, achieving real-time. Besides, the total memory usage is 7846 MB.
Inversion overhead. The time overhead of inversion depends on two factors: the number of iterations needed for convergence and the time taken for each iteration. Among them, the time taken for each iteration is 150 ms. For the number of iterations, there is a significant difference between different frames. For the first frame of a new video (or new scene), the inversion takes thousands of iterations to converge since it starts from scratch. For subsequent frames, thanks to the continuous prompt space §3.3 providing sufficient scene priors, the inversion can converge in only a few hundred iterations. We present an example in Figure 18. It can be seen that the second frame requires only 120 iterations to achieve a visual quality comparable to that of 2000 iterations for the first frame. In practice, we set the number of iterations for the first frame to 10,000, and for subsequent frames to 500. As for memory usage, inversion occupies a total of 10 GB.
@@ 6 Limitation
The time overhead of prompt fitting. Although Promptus can achieve real-time video generation at the receiver, prompt fitting cannot be performed in real-time at the sender. This is because prompt fitting requires iterative gradient descent. Although a single iteration is fast, the total time overhead is high due to the large number of iterations required for convergence (such as 500 iterations). As a result, Promptus can currently only be used for Video on Demand and is not applicable for Live Video or Real-Time Communication. To address this, using more efficient gradient descent algorithms to reduce the number of iterations required for convergence is expected to accelerate prompt fitting to real-time.
The latency of prompt interpolation. At the receiver side, Promptus obtains the prompts of intermediate frames through prompt interpolation of keyframes. This means that if an intermediate frame needs to be generated and played, it is necessary to wait until the subsequent keyframe is received, which introduces additional latency. Although this latency has little impact on Video on Demand, as DASH (sodagar2011mpeg, ) provides ample buffering to cover the interpolation latency. However, it is not suitable for latency-sensitive applications, such as WebRTC-based (razor, ) video conferencing and cloud gaming wu2023zgaming , where buffer sizes are small. To address this issue, designing keyframe extrapolation algorithms to replace interpolation is a future research direction.
Non-uniform keyframes. According to §4.1, Promptus sends keyframes uniformly based on the keyframe interval. However, different segments of a video often have different rates of change. For rapidly changing segments, the distance between keyframes in the prompt space is large, resulting in a poor approximation of intermediate frames using linear interpolation. A solution is to reduce the keyframe interval and send keyframes more densely. But there also exist smoothly changing segments in the video, where densely sending keyframes brings little improvement to quality, thus wasting bandwidth. In summary, sending keyframes uniformly in this paper is inefficient. In the future, adaptive keyframe needs to be designed to transmit densely in rapidly changing segments and sparsely in smooth segments.
Performance on objective metrics. Although Promptus achieves significant improvements in perceptual quality, it does not perform equally well on some objective metrics (such as PSNR, SSIM), as shown in Table 2. This is because Promptus is more sensitive to sharp textures like all AIGC models, while PSNR and SSIM are more sensitive to smooth textures with large areas. As a result, the images generated by Promptus have sharper details with better perceptual quality, while the baselines are more blurred but with higher PSNR. Sometimes, the more details presented by Promptus, the lower the PSNR, while the more blurred the baselines are, the higher the PSNR can be. This phenomenon is also illustrated in the subjective examples in Figure 12.
Table 2: Comparison of PSNR, SSIM, and LPIPS metrics on the UVG (mercat2020uvg, ) at a bitrate of 225 kbps.
H.265 VAE NAS VQGAN Ours
PSNR $\uparrow$ 32.685 28.008 33.466 25.740 30.378
SSIM $\uparrow$ 0.892 0.646 0.780 0.648 0.702
LPIPS $\downarrow$ 0.386 0.382 0.357 0.335 0.244
Although PSNR and SSIM are widely used in the video compression task. However, in this paper, Promptus is more oriented towards the semantic communication task. The main goal is to keep the semantic information correct at low bitrates. LPIPS considers perceptual quality, making it more suitable for evaluating semantic distortion. After all, Promptus is a novel paradigm for high-fidelity semantic communications.
@@ 7 Conclusion
In this paper, we propose Promptus, a novel system that replaces video streaming with prompt streaming by representing video frames as Stable Diffusion prompts. To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed. To achieve adaptive bitrate, a low-rank decomposition-based bitrate control algorithm is introduced. For inter-frame compression, an interpolation-aware fitting algorithm is proposed. Evaluations across various video domains and real network traces demonstrate that, compared to H.265, Promptus can achieve more than a 4x bandwidth reduction while preserving the same perceptual quality. Promptus extends the boundaries of AIGC to video streaming, offering a semantic video communication paradigm.
Promptus is now open-source, including real-time demos, generation engines, pre-fitted prompts, and comprehensive documentation: https://github.com/JiangkaiWu/Promptus.